# Neural-Driven Image Editing

### v2

@@ 5.4 Ablation Studies on Model Architecture
Finding 1: Each component in the LoongX architecture contributes uniquely, and their joint pretraining through multimodal alignment maximizes their collective strengths, leading to improved editing accuracy across metrics.
The ablation study in Table 2 is conducted using the setting of fusing all signals and speech, exploring the impact of each proposed module. CS3 encoder enhances the completeness and smoothness of features via extracted features, reducing pixel-level error (L2 decreased by $5\%$), and DGF primarily improves semantic alignment with textual instructions (CLIP-T improvement: $3.5\%$). Supplemented by pre-training, LoongX reaches optimal performance, indicating the important role of robust multimodal alignment and sementic representation learning in maximizing editing performance.
@@
5.5 Qualitative Analysis and Limitations
Finding 2: LoongX demonstrates image editing capabilities from neural signals, though performance may vary slightly with complex or ambiguous intentions.
Qualitative results shown in Fig. 7 illustrate that neural-signal-driven editing effectively handles visual and structural modifications, such as background replacing and global adjustments. However, the fused neural-language method better captures nuanced instructions involving abstract semantics (e.g., “modify the text information"). Despite significant advancements, entity consistency (e.g., the style of the little girl in Fig. 7(b)) remains a limitation of current editing models. Moreover, highly abstract or ambiguous instructions occasionally still pose challenges (e.g., “winged white animal in Fig 11" and several more failure cases shown in Fig. 14), indicating areas where further refinement of entity interpretation and disambiguation from neural data is necessary.
@@ 6 Conclusion
We presented LoongX, a novel framework for hands-free image editing by conditioning diffusion models on multimodal neural signals, achieving performance comparable to or superior to traditional text-driven baselines. Looking forward, the portability of our wireless setup opens exciting possibilities for real-world applications in immersive environments. Future directions include integrating LoongX with VR/XR to support intuitive cognitive interaction and aligning neural representations with world models for mind-driven control in virtual environments.
Appendix A Technical Appendix and Supplementary Information
@@ A.1 Preliminary: Analysis of Brain Region Function and Mechanism
EEG Ch 0 (Pz) Parietal cortex Spatial attention, visuomotor integration Focuses on specific image areas, targets object localization [58]
@@
EEG Ch 1 (Fp2) Prefrontal cortex Emotion regulation, motivational drive Generates and regulates intentional editing actions [59]
@@
EEG Ch 2 (Fpz) Frontopolar cortex Attention control, task initiation Triggers editing intention, starts/stops editing operations [60, 61]
@@
EEG Ch 3 (Oz) Occipital cortex Visual perception, image processing Perceives visual changes, evaluates whether edits meet expectations [62]
@@
fNIRS (aMPFC) Anterior medial prefrontal cortex High-level cognition, motivation, emotional valence Indicates editing intent intensity, emotional confidence, and mental workload [63]
@@
PPG (aMPFC) Cardiovascular or autonomic system Heart rate variability, arousal, stress Monitors cognitive stress or emotional arousal during editing  [31, 32]
@@
To enable hands-free image editing, we integrate multimodal neural and physiological signals to decode user intent and cognitive state in real time. The functional roles of multimodal neural signals are provided in Table.3. Specifically, midline parietal EEG (Pz) reflects spatial attention and visuomotor integration, supporting the allocation of attention to target areas and coordination of motor plans during editing tasks [58]. Right prefrontal EEG (Fp2) is linked to emotional regulation and motivational drive, aligning with findings from frontal alpha asymmetry studies [59]. Frontopolar EEG (Fz/Fpz) tracks attention control and task initiation, facilitating the onset and modulation of editing operations [60, 61]. Occipital EEG (Oz) encodes visual perception and image processing load, enabling evaluation of visual changes and edit quality [62]. fNIRS over anterior medial prefrontal cortex (aMPFC) measures cognitive load, motivation, and emotional valence through hemodynamic activity [63]. PPG signals from the same region capture heart rate variability and autonomic arousal, which reflect user stress and engagement [31, 32]. This multimodal mapping (Fig. 8 (b)) enables the system to continuously adapt to users’ cognitive focus, affective state, and mental workload, thereby supporting a more responsive and intuitive editing experience.
A.2 Subject information
[…]
This ratio has been widely used in cognitive neuroscience research as a neurophysiological index of attentional control and mental workload. For instance, Raufi [64] demonstrated that the alpha-to-theta and theta-to-alpha band ratios are reliable indicators of self-reported mental workload levels in EEG-based studies. The attention scores and the neural-signal based image editing errors of each individual subject are summarized in Table.4.
Table 4: Attention and neural-signal based image editing scores of individual subject
Subject Gender Age #Samples Attention L1 L2 CLIP-I DINO CLIP-T 1 Female 25 2003 0.0887 0.2657 0.1109 0.6370 0.4890 0.2196 2 Female 29 2000 0.0817 0.2416 0.0950 0.6575 0.5021 0.2249 3 Female 26 2001 0.1340 0.2448 0.0963 0.6660 0.4878 0.2337 4 Female 22 1999 0.0739 0.2533 0.1005 0.6394 0.4606 0.2270 5 Female 28 1992 0.1218 0.2552 0.1031 0.6144 0.4157 0.2260 6 Female 29 1964 0.0822 0.2511 0.1000 0.6449 0.4564 0.2213 7 Male 22 1988 0.0851 0.2711 0.1160 0.6515 0.4634 0.2234 8 Male 23 1993 0.1105 0.2528 0.1017 0.6638 0.4833 0.2242 9 Male 22 1988 0.1500 0.2497 0.0998 0.6355 0.4571 0.2212 10 Male 24 2000 0.1298 0.2657 0.1144 0.6194 0.4240 0.2220 11 Male 24 2000 0.0954 0.2744 0.1151 0.6386 0.4339 0.2299 12 Male 22 2000 0.0971 0.2551 0.1034 0.6213 0.4323 0.2250
A.3 Preliminary Analysis: Editing Type Classification Experiment

### v3

@@ 5.5 Ablation Studies on Model Architecture
Each component in the LoongX architecture contributes uniquely, and their composition (especially with pretraining) maximizes the performance potential, leading to improved editing accuracy across metrics. The ablation study in Table 2 is conducted on the fused neural signals and speech to evaluate the impact of each proposed module. It is found that CS3 encoder enhances feature completeness and smoothness, leading to a 5% reduction in L2 error. The DGF module improves semantic alignment with textual instructions, yielding a 3.5% increase in CLIP-T. Supplemented by pre-training, LoongX reaches optimal performance, indicating the important role of robust and representation learning and multimodal alignment in maximizing editing performance.
@@
5.6 Qualitative Analysis and Discussion
Qualitative examples confirm LoongX’s intuitive editing capabilities, though performance may vary slightly in scenarios with complex or ambiguous intentions. Qualitative results presented in Fig. 7 demonstrate that neural-driven LoongX can successfully achieve various visual and structural modifications, such as background replacement and global adjustments. However, the fused neural-language method better captures nuanced instructions involving abstract semantics (e.g., “modify the text information" in Fig. 7(d)).
Limitations and future work.
It is noted that while the neural signals and combined methods perform better in multiple tasks such as object manipulation (e.g., letting the cat look down in Fig. 7(a)), the text-based method handles spatial manipulation more effectively (e.g., “place the cat above" in Fig. 7(a)). Despite significant advancements with multimodal fusion, entity consistency (e.g., the style of the little girl in Fig. 7(b)) remains a challenge, which is limited by the capabilities of backbone image editing models at the time this work was mainly completed. Moreover, highly abstract or ambiguous instructions occasionally still pose challenges (e.g., “winged white animal in Fig. 11(f)"). Several more failure cases are shown in Fig. 14, indicating areas where further refinement in entity interpretation and neural-data-based disambiguation remains necessary.
Overall, the experiments validate LoongX’s efficacy as a robust, intuitive interface leveraging neural signals for image editing. Multimodal integration, particularly neural and linguistic fusion, emerges as essential for capturing comprehensive user intent. These findings advocate for future research focusing on improving semantic interpretation fidelity from neural signals and exploring adaptive methodologies for further enhancing accessibility and precision in cognitive-driven creative technologies. Incremental fine-tuning for new unseen users is also worth exploration in practice.
@@ 6 Conclusion
We presented LoongX, a novel framework for hands-free image editing by conditioning diffusion models on multimodal neural signals, achieving performance comparable to or superior to traditional text-driven baselines. Looking ahead, the portability of our wireless setup opens exciting possibilities for real-world applications in immersive virtual environments. Future directions include integrating LoongX with VR/XR to support intuitive cognitive interaction and aligning neural representations with emerging world models Mao et al. (2025b); Ni et al. (2025c) to project human intention into an interactive synthetic world, enabling mind-driven control in virtual realities.
Acknowledgments and Disclosure of Funding
We sincerely thank Prof. Tat-Seng Chua (National University of Singapore) for his valuable comments and suggestions during the drafting of this work. We also thank Dr. Zhiwei Tang (Alibaba) for his constructive guidance during the rebuttal. This work is sponsored by the NUS Startup Grant (Presidential Young Professorship), Singapore MOE Tier-1 Grant, ByteDance Grant, NUS ARTIC Grant, Apple Grant, Alibaba Grant, and Google Grant for TPU usage, and also supported by the STI2030-Major Projects (No. 2021ZD0200401), Zhejiang Province Key R&D Programs (No. 2024C03001, 2024C03007, 2024C01031, 2025C01137, 2025C02165), Zhejiang Province High-Level Talent Special Support Plan (No. 2022R52042), and the Fundamental Research Funds for the Central Universities (No. 2025ZFJH01).
Appendix A Supplementary Key Information
@@ A.1 Preliminary: Analysis of Brain Region Function and Mechanism
EEG Ch 0 (Pz) Parietal cortex Spatial attention, visuomotor integration Focuses on specific image areas, targets object localization (Praamstra et al., 2005)
@@
EEG Ch 1 (Fp2) Prefrontal cortex Emotion regulation, motivational drive Generates and regulates intentional editing actions (Sengupta et al., 2024)
@@
EEG Ch 2 (Fpz) Frontopolar cortex Attention control, task initiation Triggers editing intention, starts/stops editing operations (Friedman and Robbins, 2022; Kim et al., 2021)
@@
EEG Ch 3 (Oz) Occipital cortex Visual perception, image processing Perceives visual changes, evaluates whether edits meet expectations (Özkara, 2019)
@@
fNIRS (aMPFC) Anterior medial prefrontal cortex High-level cognition, motivation, emotional valence Indicates editing intent intensity, emotional confidence, and mental workload (Agrawal et al., 2025)
@@
PPG (aMPFC) Cardiovascular or autonomic system Heart rate variability, arousal, stress Monitors cognitive stress or emotional arousal during editing  (Arsalan and Majid, 2021; Tazarv et al., 2021)
@@
To enable hands-free image editing, we integrate multimodal neural and physiological signals to decode user intent and cognitive state in real time. The functional roles of multimodal neural signals are provided in Table.3. Specifically, midline parietal EEG (Pz) reflects spatial attention and visuomotor integration, supporting the allocation of attention to target areas and coordination of motor plans during editing tasks (Praamstra et al., 2005). Right prefrontal EEG (Fp2) is linked to emotional regulation and motivational drive, aligning with findings from frontal alpha asymmetry studies (Sengupta et al., 2024). Frontopolar EEG (Fz/Fpz) tracks attention control and task initiation, facilitating the onset and modulation of editing operations (Friedman and Robbins, 2022; Kim et al., 2021). Occipital EEG (Oz) encodes visual perception and image processing load, enabling evaluation of visual changes and edit quality (Özkara, 2019). fNIRS over anterior medial prefrontal cortex (aMPFC) measures cognitive load, motivation, and emotional valence through hemodynamic activity (Agrawal et al., 2025). PPG signals from the same region capture heart rate variability and autonomic arousal, which reflect user stress and engagement (Arsalan and Majid, 2021; Tazarv et al., 2021). This multimodal mapping (Fig. 8(b)) enables the system to continuously adapt to users’ cognitive focus, affective state, and mental workload, thereby supporting a more responsive and intuitive editing experience.
A.2 Preliminary Analysis: Editing Type Classification Experiment